OpenStudy (anonymous):

The tickets for a high school basketball game are $2.50 for students and $4.00 for adults. At Tuesday night's game, 95 tickets were sold for a total of $290.00. How many student tickets were sold? 60 58 35 33

OpenStudy (saifoo.khan):

x + y = 95 2.5x + 4y = 290

OpenStudy (saifoo.khan):

Who want to solve further?

OpenStudy (saifoo.khan):

60 for children. 35 for adults.

OpenStudy (anonymous):

S students A adults 2.5*S + 4A = $290 S + A = 95 multiply S + A = 95 by -4 so you get these 2 eq's 2.5*S + 4A = $290 -4*S - 4A = -380 Add them -1.5*S = -90 So, S = 60 60 + A = 95, so adults = 35